OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (aum):

-5 is negative. |x+1| is always positive or zero. Can the product of those two be a positive number for any value of x?

OpenStudy (anonymous):

yes? I think but am also confused why is it always positive or zero?

OpenStudy (aum):

That is what the absolute bar does. |-8| = 8 So the absolute operator turns negative values to positive values.

OpenStudy (aum):

−5|x + 1| = 10 divide both sides by -5 |x+1| = -2

OpenStudy (anonymous):

wait it cant be -2

OpenStudy (aum):

Yes |x+1| can never be a negative number. So |x+1| = -2 has NO solution. There is no value of x, positive, zero or negative that would make |x+1| equal negative two. So there is no solution to this problem.

OpenStudy (anonymous):

oh srry I see now thanks again